# Computation of Robust Option Prices via Structured Multi-Marginal Martingale Optimal Transport

### v1

@@ 4 Solving the structured problem via regularisation
We have now arrived at a version of the problem that can be approached computationally. We therefore proceed by deriving a generalisation of the algorithm described in Section 2.3, modified to also take the martingale constraint into account. As in Section 2.3, this is done by adding a regularising entropy term $D(Q):=\sum_{\boldsymbol{\omega}\in\boldsymbol{\Omega}}(Q(\boldsymbol{\omega})\log Q(\boldsymbol{\omega})\allowbreak-Q(\boldsymbol{\omega})+1)$, scaled by some small number $\varepsilon>0$, to the objective function of problem (20). The entropy regularised version of problem (13) written on tensor form thus becomes
@@
$\underset{Q\in\mathbb{R}_{+}^{(n_{S}n_{X})^{T+1}}}{\min}\quad$ $\langle C,Q\rangle+\varepsilon D(Q)$ (22a)
subject to $\left(P_{t-1,t}(Q)\odot\Delta\right)\mathbf{1}_{n_{S}n_{X}}=\mathbf{0}_{n_{S}n_{X}},$ $t\in[T]\backslash 0$ (22b)
$P_{t}^{S}(Q)=\upmu_{t},$ $t\in\mathcal{T},$ (22c)
@@
where we recall that the tensor $C$ is of the form $C(\boldsymbol{\omega})=\sum_{[T]\backslash 0}C_{t}(\omega_{t-1},\omega_{t})$, $\boldsymbol{\omega}\in\boldsymbol{\Omega}$, for matrices $C_{t}\in\mathbb{R}^{n_{S}n_{X}\times n_{S}n_{X}}$. Note that the above problem is no longer linear, but rather strictly convex, due to the entropy term $D(Q)$ being strictly convex in $Q$. Since it is a convex problem whose feasible set has non-empty relative interior, one can show (see, for example, [47]) that strong duality holds. We can therefore solve (22) by considering the dual problem. We will see that the optimality conditions for the dual variables correspond to relatively simple equations, where explicit expressions can be obtained for the dual variables corresponding to marginal constraints, while for the dual variables corresponding to martingale constraints it turns out that $T$ sets of equations must be solved numerically. This is a generalisation of the strategy used by [18] for the bi-marginal MOT problem. We then exploit the structure inherent in the problem to simplify the equations obtained; the resulting algorithm is a multi-marginal version of Sinkhorn’s algorithm where some of the variables are updated by using Newton’s method.
4.1 Lagrangian relaxation and coordinate dual ascent
In order to derive the dual of problem (22), we start by forming the corresponding Lagrangian function. It reads
$\mathcal{L}(Q,\lambda,\gamma)=\langle C,Q\rangle+\varepsilon D(Q)+\sum_{t\in\mathcal{T}}\big{[}\lambda_{t}^{\top}\left(\upmu_{t}-P^{S}_{t}(Q)\right)\big{]}-\sum_{t\in[T]\backslash 0}\big{[}\gamma_{t-1}^{\top}\left(P_{t-1,t}(Q)\odot\Delta\right)\mathbf{1}_{n_{S}n_{X}}\big{]},$
where $\lambda=\{\lambda_{t}\}_{t\in\mathcal{T}}$ are Lagrangian multipliers (dual variables) corresponding to the marginal constraints (22c) and $\gamma=\{\gamma_{t}\}_{t=0}^{T-1}$ are Lagrangian multipliers corresponding to the martingale constraints (22b). Note that each $\lambda_{t}$ is a vector of length $n_{S}$, while each $\gamma_{t}$ is a vector of length $n_{S}n_{X}$. The dual of problem (22) is then
$\underset{\lambda,\gamma}{\max}\quad\varphi(\lambda,\gamma),$
where the dual functional is defined as $\varphi(\lambda,\gamma):=\min\{\mathcal{L}(Q,\lambda,\gamma):Q\in\mathbb{R}_{+}^{(n_{S}n_{X})^{T+1}}\}$. We proceed by finding the minimising $Q_{\lambda,\gamma}$ for given dual variables $\lambda,\gamma$. Since the Lagrangian function $\mathcal{L}$ is strictly convex in $Q$ and the entropy term $D(Q)$ prevents optimal solutions at the boundary, the minimising transport plan satisfies $\partial_{Q(\boldsymbol{\omega})}\mathcal{L}(Q,\lambda,\gamma)=0$ for $\boldsymbol{\omega}\in\boldsymbol{\Omega}$. By differentiating the Lagrangian function we obtain that for any $\boldsymbol{\omega}\in\boldsymbol{\Omega}$
$\partial_{Q(\boldsymbol{\omega})}\mathcal{L}(Q,\lambda,\gamma)=$ $\varepsilon\log\left(Q(\boldsymbol{\omega})\right)-\sum_{t\in\mathcal{T}}\lambda_{t}(\omega_{t}^{S})+\sum_{t\in[T]\backslash 0}\left(C_{t}(\omega_{t-1},\omega_{t})-\gamma_{t-1}(\omega_{t-1})\Delta(\omega_{t-1},\omega_{t})\right)$
which yields a minimising transport plan
$Q_{\lambda,\gamma}(\boldsymbol{\omega})=\Bigg{(}\prod_{t\in\mathcal{T}}e^{\lambda_{t}(\omega_{t}^{S})/\varepsilon}\Bigg{)}\Bigg{(}\prod_{t\in[T]\backslash 0}e^{-C_{t}(\omega_{t-1},\omega_{t})/\varepsilon}e^{\gamma_{t-1}(\omega_{t-1})\Delta(\omega_{t-1},\omega_{t})/\varepsilon}\Bigg{)},\quad\boldsymbol{\omega}\in\boldsymbol{\Omega},$ (23)
for which the first term of the scaled entropy $\varepsilon D(Q_{\lambda,\gamma})$ equals
$\varepsilon\sum_{\boldsymbol{\omega}\in\boldsymbol{\Omega}}$ $Q_{\lambda,\gamma}(\boldsymbol{\omega})\log\left(Q_{\lambda,\gamma}(\boldsymbol{\omega})\right)$
$=\sum_{\boldsymbol{\omega}\in\boldsymbol{\Omega}}Q_{\lambda,\gamma}(\boldsymbol{\omega})\Bigg{(}\sum_{t\in\mathcal{T}}\lambda_{t}(\omega_{t}^{S})-\sum_{t\in[T]\backslash 0}\left(C_{t}(\omega_{t-1},\omega_{t})-\gamma_{t-1}(\omega_{t-1})\Delta(\omega_{t-1},\omega_{t})\right)\Bigg{)}$
$=\sum_{t\in\mathcal{T}}\lambda_{t}^{\top}P_{t}^{S}(Q_{\lambda,\gamma})-\langle C,Q_{\lambda,\gamma}\rangle+\sum_{t\in[T]\backslash 0}\gamma_{t-1}^{\top}\left(P_{t-1,t}(Q_{\lambda,\gamma})\odot\Delta\right)\mathbf{1}_{n_{S}n_{X}}.$
Inserting $Q_{\lambda,\gamma}$ into $\mathcal{L}(\cdot,\lambda,\gamma)$ thus yields that the objective of the dual of problem (22) becomes
$\varphi$ $(\lambda,\gamma)$
$=\sum_{t\in\mathcal{T}}\lambda_{t}^{\top}\upmu_{t}+\varepsilon\sum_{\boldsymbol{\omega}\in\boldsymbol{\Omega}}\Bigg{[}1-\Bigg{(}\prod_{t\in\mathcal{T}}e^{\lambda_{t}(\omega_{t}^{S})/\varepsilon}\Bigg{)}\Bigg{(}\prod_{t\in[T]\backslash 0}e^{-C_{t}(\omega_{t-1},\omega_{t})/\varepsilon}e^{\gamma_{t-1}(\omega_{t-1})\Delta(\omega_{t-1},\omega_{t})/\varepsilon}\Bigg{)}\Bigg{]}.$
At this point, we note that $\varphi$ can be expressed in a more compact format — a format that reduces the number of direct applications of the exponential function later on — by reparameterising the dual variables. Therefore, let $u_{t}$ for $t\in\mathcal{T}$ be a family of vectors in $\mathbb{R}^{n_{S}n_{X}}$ given by
$u_{t}(\omega_{t}^{S}):=e^{\lambda_{t}(\omega_{t}^{S})/\varepsilon},\quad\omega_{t}^{S}=1,\dots,n_{S},$ (24)
and let $K_{t},G_{t}$, for $t\in[T]\backslash 0$, be two families of matrices in $\mathbb{R}^{n_{S}n_{X}\times n_{S}n_{X}}$, defined by
$K_{t}(\omega_{t-1},\omega_{t}):=e^{-C_{t}(\omega_{t-1},\omega_{t})/\varepsilon},\quad G_{t}(\omega_{t-1},\omega_{t}):=e^{\gamma_{t-1}(\omega_{t-1})\Delta(\omega_{t-1},\omega_{t})/\varepsilon},$ (25)
for $\omega_{t-1},\omega_{t}\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}$. Then $u=\{u_{t}\}_{t\in\mathcal{T}}$ are variables associated with the marginal constraints (22c), while $G_{t}$ is a matrix representing the $t^{\text{th}}$ martingale constraint (22b). It connects the $(t-1)^{\text{th}}$ and $t^{\text{th}}$ marginals to each other. The matrix $K_{t}$ represents the cost associated with moving from the $(t-1)^{\text{th}}$ to the $t^{\text{th}}$ marginal. Let $K,U_{\lambda},G_{\gamma}\in\mathbb{R}^{(n_{S}n_{X})^{T+1}}$ be tensors defined by
@@ 4.1 Lagrangian relaxation and coordinate dual ascent
$K(\boldsymbol{\omega}):=\prod_{t\in[T]\backslash 0}K_{t}(\omega_{t-1},\omega_{t}),\quad U_{\lambda}(\boldsymbol{\omega}):=\prod_{t\in\mathcal{T}}u_{t}(\omega_{t}^{S}),\quad G_{\gamma}(\boldsymbol{\omega}):=\prod_{t\in[T]\backslash 0}G_{t}(\omega_{t-1},\omega_{t}),$ (26)
for $\boldsymbol{\omega}\in\boldsymbol{\Omega}$. Also define
$U_{\lambda}^{-r}(\boldsymbol{\omega}):=\prod_{t\in\mathcal{T}\backslash r}u_{t}(\omega_{t}^{S}),\quad r\in\mathcal{T},\quad G_{\gamma}^{-q}(\boldsymbol{\omega}):=\prod_{t\in[T]\backslash\{0,q\}}G_{t}(\omega_{t-1},\omega_{t}),\quad q\in[T]\backslash 0,\quad\boldsymbol{\omega}\in\boldsymbol{\Omega}.$
In this notation, the expression for $Q_{\lambda,\gamma}$ in equation 23 becomes
$Q_{\lambda,\gamma}=K\odot U_{\lambda}\odot G_{\gamma}.$ (27)
This is the minimising primal variable $Q$, given dual variables $\lambda,\gamma$. The dual functional can be written as $\varphi(\lambda,\gamma)=\sum_{t\in\mathcal{T}}\lambda_{j}^{\top}\upmu_{j}-\varepsilon\langle K,U_{\lambda}\odot G_{\gamma}\rangle+\varepsilon(n_{S}n_{X})^{T+1}$ using the new notation; the dual problem thus reads
@@
$\max_{\lambda,\gamma}\quad\sum_{t\in\mathcal{T}}\lambda_{j}^{\top}\upmu_{j}-\varepsilon\langle K,U_{\lambda}\odot G_{\gamma}\rangle+\varepsilon(n_{S}n_{X})^{T+1}.$ (28)
It remains to find $\lambda,\gamma$ such that the concave functional $\varphi$ is maximised — that is, that satisfies the following set of equations
$\partial_{\lambda_{t}(i)}\varphi(\lambda,\gamma)=0,\quad t\in\mathcal{T},\quad i\in\{1,\dots,n_{S}\}$ (29a)
$\partial_{\gamma_{t}(i)}\varphi(\lambda,\gamma)=0,\quad t\in[T-1],\quad i\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}.$ (29b)
Once they are known, they can be used to recover the optimal transport plan $Q_{\lambda,\gamma}$ via equations (24) – (27).
$\partial_{\lambda_{t}(i)}\varphi(\lambda,\gamma)=$ $\upmu_{t}(i)-\sum_{\begin{subarray}{c}\boldsymbol{\omega}\in\boldsymbol{\Omega}:\\ \omega_{t}^{S}=i\end{subarray}}K(\boldsymbol{\omega})G_{\gamma}(\boldsymbol{\omega})\prod_{r\in\mathcal{T}}e^{\lambda_{r}(\omega_{r}^{S})/\varepsilon},\quad t\in\mathcal{T},\quad i\in\{1,\dots,n_{S}\}.\quad$
Solving for $u_{t}=\exp(\lambda_{t}/\varepsilon)$ thus yields
$u_{t}=\upmu_{t}\oslash P_{t}^{S}(K\odot U_{\lambda}^{-t}\odot G_{\gamma}),\quad t\in\mathcal{T}.$ (30)
This is the equation characterising the representation $u$ of the optimal dual variables corresponding to marginal constraints in problem (22). Note that the right-hand side of equation 30 can be written as $(\upmu_{t}\odot u_{t})\oslash P^{S}_{t}(K\odot U_{\lambda}\odot G_{\gamma})$. Equation 30 thus requires that $u_{t}$ should be chosen such that the corresponding $Q_{\lambda,\gamma}$ satisfies $P^{S}_{t}(Q_{\lambda,\gamma})=\upmu_{t}$, which we recognise as the constraint (22c).
Moving on to the dual variables corresponding to martingale constraints in the primal problem, we differentiate and get
$\partial_{\gamma_{t}(j)}\varphi(\lambda,\gamma)=-\sum_{\begin{subarray}{c}\boldsymbol{\omega}\in\boldsymbol{\Omega}:\\ \omega_{t}=j\end{subarray}}\Delta(j,\omega_{t+1})K(\boldsymbol{\omega})U_{\lambda}(\boldsymbol{\omega})\prod_{t\in[T]\backslash 0}e^{\gamma_{r-1}(\omega_{r-1})\Delta(\omega_{r-1},\omega_{r})/\varepsilon},$
for $t\in[T-1]$ and $j\in\{1,..,n_{S}\}\times\{1,\dots,n_{X}\}$. When solving for $\gamma$ such that $\partial_{\gamma_{t}(j)}\varphi(\lambda,\gamma)=0$ for all $t$ and $j$ as given above, we do not get any explicit formulas. Instead we obtain for every $t\in[T-1]$, $n_{S}n_{X}$ equations in $\gamma_{t}(j)$, $j\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}$. They can be written on vector form as
$\left(P_{t,t+1}(K\odot U_{\lambda}\odot G_{\gamma})\odot\Delta\right)\mathbf{1}_{n_{S}n_{X}}=\mathbf{0}_{n_{S}n_{X}},\quad t\in[T-1],$ (31)
where the above should be solved with respect to $\gamma_{t}$, noting that $G_{\gamma}=\exp(\operatorname{diag}(\gamma_{t})\Delta/\varepsilon)\odot G_{\gamma}^{-(t+1)}$. We recognise equation 31 as constraint (22b). We emphasise that there are no cross-dependencies between the elements of $\gamma_{t}$ in equation 31; this will turn out to be useful when deriving a Newton scheme later on.
@@
Initialise: $u_{t}\leftarrow\mathbf{1}_{n_{S}}$ for $t\in[T]$
Initialise: $\gamma_{t}\leftarrow\mathbf{1}_{n_{S}n_{X}}$ for $t\in[T-1]$
@@
         $u_{t}\leftarrow\upmu_{t}\oslash P_{t}^{S}\left(K\odot U_{\lambda}^{-t}\odot G_{\gamma}\right)$
@@
         find $\hat{\gamma}_{t}$ such that $(P_{t,t+1}(K\odot U_{\lambda}\odot G_{\gamma})\odot\Delta)\boldsymbol{1}_{n_{S}n_{X}}=\boldsymbol{0}_{n_{S}n_{X}}$
@@
$Q_{\lambda,\gamma}\leftarrow K\odot U_{\lambda}\odot G_{\gamma}$
return $Q_{\lambda,\gamma}$

### v2

@@ 4 Solving the structured problem via regularisation
We have now arrived at a version of the problem that can be approached computationally and we therefore proceed by deriving a generalisation of the algorithm described in Section 2.3, modified to also take the martingale constraint into account. As in Section 2.3, this is done by adding a regularising entropy term [^6] Our definition of the discrete entropy $D(\mathbf{Q})$ corresponds to using the uniform measure as reference measure in the Kullback-Leibler divergence. Other choices of such measures could be made though — since the reference measure can be interpreted as the distribution of some reference stochastic process, this would be natural in a context where such a process is available. Martingale couplings that minimise the Kullback-Leibler divergence with respect to this distribution would then be favored. However, using another reference measure would affect the objective function of the regularised problem (20) only slightly when the regularisation parameter $\varepsilon$ is small. Indeed, let the tensor associated with a general reference measure in the case of a finite state space be denoted by $\mathbf{R}$. Then it is straightforward to show that the new objective function of the regularised problem can be written $\langle\mathbf{C}-\varepsilon\log\mathbf{R},\mathbf{Q}\rangle+\varepsilon D(\mathbf{Q})$. $D(\mathbf{Q})=\langle\log(\mathbf{Q})-\mathbf{1}_{n_{0}\times\dots\times n_{T}},\mathbf{Q}\rangle$, scaled by some small number $\varepsilon>0$, to the objective function of problem (19). The entropy regularised version of problem (13) written on tensor form thus becomes
@@
$\underset{\mathbf{Q}\in\mathbb{R}_{+}^{n_{0}\times\dots\times n_{T}}}{\min}\quad$ $\langle\mathbf{C},\mathbf{Q}\rangle+\varepsilon D(\mathbf{Q})$ (20a)
subject to $\left(P_{t-1,t}(\mathbf{Q})\odot\Delta_{t}\right)1_{n_{t}}=0_{n_{t-1}},$ $t\in[T]\backslash 0$ (20b)
$P_{t}^{\mathcal{S}}(\mathbf{Q})=m_{t},$ $t\in\mathcal{T},$ (20c)
where we recall that the tensor $\mathbf{C}$ is of the form $\mathbf{C}(i_{0},\dots,i_{T})=\sum_{[T]\backslash 0}C_{t}(i_{t-1},i_{t})$, $(i_{0},\dots,i_{T})\in\mathcal{I}$, for matrices $C_{t}\in\mathbb{R}^{n_{t-1}\times n_{t}}$. Note that a minimum exists for the above problem, since the objective function is continuous in $\mathbf{Q}$ and the feasible set is compact [^7] The marginal constraints ensures that the feasible set of problem (20) is bounded. Compactness then follows from the continuity of the constraints (20b) and (20c)., and that it is unique since the problem is strictly convex. We now show convergence of optimal solutions of the regularised problem as the regularisation parameter vanishes. The proof can be found in Appendix: Additional proofs. It follows that a solution of problem (20) serves as an approximate solution of problem (19) when $\varepsilon$ is small.
Proposition 4.1.
Let $(\varepsilon_{k})_{k}$ be a decreasing sequence of positive regularisation parameters such that $\lim_{k\rightarrow\infty}\varepsilon_{k}=0$ and let $\mathbf{Q}_{k}$ denote the optimal solution of the regularised problem (20) with $\varepsilon=\varepsilon_{k}$. Then the sequence $(\mathbf{Q}_{k})_{k}$ of minimisers has at least one convergent subsequence and every limit point of $(\mathbf{Q}_{k})_{k}$ is a minimiser of the linear programming problem (19). Moreover, the value of the regularised problem (20) converges to the value of problem (19) as the regularisation parameter vanishes.
@@
Next, we will see that strong duality holds for problem (20), and that under certain additional assumptions on the given marginals a dual maximiser exists and problem (20) can then be solved by considering the dual problem. We will also see that the optimality conditions for the dual variables correspond to relatively simple equations, where explicit expressions can be obtained for the dual variables corresponding to marginal constraints, while for the dual variables corresponding to martingale constraints it turns out that $T$ sets of equations must be solved numerically. This is a generalisation of the strategy used by [21] for the bi-marginal MOT problem. We then exploit the structure inherent in the problem to simplify the equations obtained; the resulting algorithm is a multi-marginal version of Sinkhorn’s algorithm where some of the variables are updated by use of Newton’s method.
4.1 Strong duality for the regularised problem
In order to provide the dual of problem (20), we introduce some new notation that allows for writing the expressions in a more compact format reducing the number of direct applications of the exponential function later on. Therefore, let $\lambda=\{\lambda_{t}\}_{t\in\mathcal{T}}$ and $\gamma=\{\gamma_{t}\}_{t\in[T-1]}$ be families of vectors with $\lambda_{t}\in\mathbb{R}^{n_{t}^{\mathcal{S}}}$ and $\gamma_{t}\in\mathbb{R}^{n_{t}}$, respectively. Given $\lambda,\gamma$ and the cost matrices $C_{t}$ for $t\in[T]\backslash 0$, define another family $u^{\lambda}=\{u_{t}^{\lambda_{t}}\}_{t\in\mathcal{T}}$ of vectors with $u_{t}^{\lambda_{t}}\in\mathbb{R}^{n_{t}^{\mathcal{S}}}$ given by
$u_{t}^{\lambda_{t}}(i_{t}^{\mathcal{S}}):=e^{\lambda_{t}(i_{t}^{\mathcal{S}})/\varepsilon},\quad i_{t}^{\mathcal{S}}\in\mathcal{I}_{t}^{\mathcal{S}}$ (21)
and let $K_{t},G^{\gamma_{t-1}}_{t}\in\mathbb{R}^{n_{t-1}\times n_{t}}$ for $t\in[T]\backslash 0$ be two families of matrices, defined by
$K_{t}(i_{t-1},i_{t}):=e^{-C_{t}(i_{t-1},i_{t})/\varepsilon},\quad G^{\gamma_{t-1}}_{t}(i_{t-1},i_{t}):=e^{\gamma_{t-1}(i_{t-1})\Delta_{t}(i_{t-1},i_{t})/\varepsilon},$ (22)
for $i_{t-1}\in\mathcal{I}_{t-1},i_{t}\in\mathcal{I}_{t}$. Let $\mathbf{K},\mathbf{U}^{\lambda},\mathbf{G}^{\gamma}\in\mathbb{R}^{n_{0}\times\dots\times n_{T}}$ be tensors defined by
$\mathbf{K}(i_{0},\dots,i_{T}):=\prod_{t\in[T]\backslash 0}K_{t}(i_{t-1},i_{t}),$ $\mathbf{U}^{\lambda}(i_{0},\dots,i_{T}):=\prod_{t\in\mathcal{T}}u_{t}^{\lambda_{t}}(i_{t}^{\mathcal{S}}),$ (23)
$\mathbf{G}^{\gamma}(i_{0},\dots,i_{T}):=\prod_{t\in[T]\backslash 0}G^{\gamma_{t-1}}_{t}(i_{t-1},i_{t}),$ $(i_{0},\dots,i_{T})\in\mathcal{I},$
and also define, for $j\in\mathcal{T}$ and $k\in[T]\backslash 0$,
@@ 4.1 Strong duality for the regularised problem
$\mathbf{U}^{\lambda}_{-j}(i_{0},\dots,i_{T}):=\prod_{t\in\mathcal{T}\backslash j}u_{t}^{\lambda_{t}}(i_{t}^{\mathcal{S}}),\quad\mathbf{G}^{\gamma}_{-k}(i_{0},\dots,i_{T}):=\prod_{t\in[T]\backslash\{0,k\}}G^{\gamma_{t-1}}_{t}(i_{t-1},i_{t}),\quad(i_{0},\dots,i_{T})\in\mathcal{I}.$
Using this notation we now state our duality result. In order to show existence of dual optimal solutions, we require that every pair $(\mu_{t_{1}},\mu_{t_{2}})$ of two consecutive given marginals with $t_{1}<t_{2}$ is irreducible according to [7, Definition A.3]; in our discrete setting this is equivalent to $|s_{t_{2}}-z1_{n^{\mathcal{S}}_{t_{2}}}|^{\top}m_{t_{2}}-|s_{t_{1}}-z1_{n^{\mathcal{S}}_{t_{1}}}|^{\top}m_{t_{1}}>0$ for $z\in\mathcal{S}_{t_{1}}\cup\mathcal{S}_{t_{2}}^{0}$, where $\mathcal{S}_{t_{2}}^{0}$ denotes $\mathcal{S}_{t_{2}}$ without its largest and smallest elments.
Theorem 4.2.
A dual of problem (20) is given by
@@
$\sup_{\lambda,\gamma}\quad\sum_{t\in\mathcal{T}}\lambda_{t}^{\top}m_{t}-\varepsilon\langle\mathbf{K},\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma}\rangle,$ (24)
where $\lambda$ and $\gamma$ are families of vectors of the above given form, and the values of the two problems are equal — that is, strong duality holds.
Moreover, suppose that the given marginals $\{\mu_{t}\}_{t\in\mathcal{T}}$ are such that there exist probability measures $\mu_{t}$ with support $\mathcal{S}_{t}$ for $t\in[T]\backslash\mathcal{T}$ such that every pair $(\mu_{t-1},\mu_{t})$ for $t\in[T]\backslash 0$ is irreducible. Then the supremum in the dual problem (24) is attained. Any dual maximiser $(\lambda^{*},\gamma^{*})$ is given as the solution of the equations
$u_{t}^{\lambda_{t}}=m_{t}\oslash P_{t}^{\mathcal{S}}(\mathbf{K}\odot\mathbf{U}^{\lambda}_{-t}\odot\mathbf{G}^{\gamma}),$ $t\in\mathcal{T},$ (25a)
$\big{(}P_{t,t+1}(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma})\odot\Delta_{t+1}\big{)}1_{n_{t+1}}=0_{n_{t}},$ $t\in[T-1],$ (25b)
and the unique martingale transport plan that minimises problem (20) is given by
$\mathbf{Q}^{\lambda^{*},\gamma^{*}}=\mathbf{K}\odot\mathbf{U}^{\lambda^{*}}\odot\mathbf{G}^{\gamma^{*}},$ (26)
for any optimal dual variables $(\lambda^{*},\gamma^{*})$. It assigns a non-zero probability mass to every trajectory that respects equation 4, that is, $\mathbf{Q}^{\lambda^{*},\gamma^{*}}(i_{0},\dots,i_{T})>0$ for $(i_{0},\dots,i_{T})\in\mathcal{I}$ such that $\mathbf{K}(i_{0},\dots,i_{T})>0$.
We refer to equation 25 as the dual optimality conditions of problem (20). In order to prove the second part of Theorem 4.2 we need the following lemma. The proofs of Theorem 4.2 and Lemma 4.3 are deferred to Appendix: Additional proofs.
Lemma 4.3.
Suppose that the grid ${}_{t\in[T]}\mathcal{S}_{t}$ and the given marginals $\{\mu_{t}\}_{t\in\mathcal{T}}$ satisfy the assumptions of the second part of Theorem 4.2. Then the domain of problem (20) has a non-empty relative interior, that is, Slater’s condition holds.
We remark that the second part of Theorem 4.2 is formulated under additional assumptions on the grid and on the given marginals compared to what have been imposed earlier in this article. We emphasise that it is always possible to construct intermediate marginals as described in Theorem 4.2 whenever each consecutive pair of marginals from the given family $\{\mu_{t}\}_{t\in\mathcal{T}}$ is irreducible and when for $t\in[T]\backslash 0$ the convex hull of the set $\mathcal{S}_{t-1}$ is a proper subset of the convex hull of $\mathcal{S}_{t}$, given that the grid is fine enough. Indeed, one can then do so by forming convex combinations of the given marginals and moving some of the probability masses. In fact, we could in theory relax the irreducibility assumption on the given marginals by noting that this would result in trajectories in the path space $\mathcal{S}_{0}\times\dots\times\mathcal{S}_{T}$ that are of zero-support for all martingale transport plans [8, pp. 3044–3045] and thus simply omit the representation of such trajectories from the problem. This would however be unpractical to implement programmatically, for what reason we have chosen the approach given here.
@@
Initialise: $u_{t}^{\lambda_{t}}\leftarrow 1_{n_{t}^{\mathcal{S}}}$ for $t\in[T]$
Initialise: $\gamma_{t}\leftarrow 1_{n_{t}}$ for $t\in[T-1]$
@@
         $u_{t}^{\lambda_{t}}\leftarrow m_{t}\oslash P_{t}^{\mathcal{S}}\left(\mathbf{K}\odot\mathbf{U}^{\lambda}_{-t}\odot\mathbf{G}^{\gamma}\right)$
@@
         find $\hat{\gamma}_{t}$ such that $(P_{t,t+1}(\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma}_{-(t+1)})\odot G_{t+1}^{\hat{\gamma}_{t}}\odot\Delta_{t+1})1_{n_{t+1}}=0_{n_{t}}$
@@
$\mathbf{Q}^{\lambda,\gamma}\leftarrow\mathbf{K}\odot\mathbf{U}^{\lambda}\odot\mathbf{G}^{\gamma}$
return $\mathbf{Q}^{\lambda,\gamma}$
@@
Remark 4.4.
We conclude this section with a note on the connection to robust hedging duality. The linear programming problem (19) has a dual
$\underset{\lambda,\gamma}{\max}\quad$ $\sum_{t\in\mathcal{T}}\lambda_{t}^{\top}m_{t}$ (27a)
subject to $\sum_{t\in[T]\backslash 0}\gamma_{t-1}(i_{t-1})\Delta_{t}(i_{t-1},i_{t})+\sum_{t\in\mathcal{T}}\lambda_{t}(i_{t}^{\mathcal{S}})\leq\mathbf{C}(i_{0},\dots,i_{T}),$ $(i_{0},\dots,i_{T})\in\mathcal{I},$ (27b)
that can be viewed as the robust subhedging problem of a derivative whose payoff is given by the tensor $\mathbf{C}$. Note that an optimal solution to problem (27) exists, since primal boundedness and feasibility implies dual feasibility and boundedness via strong linear programming duality. The dual variables $(\lambda,\gamma)$ then has an interpretation as a hedging strategy, where the variables $\lambda$ corresponds to static trading in European options and the variables $\gamma$ corresponds to dynamic trading in the underlying: the element $\lambda_{t}(j)$, $j\in\mathcal{I}_{t}^{\mathcal{S}}$, provides the payoff of European options expiring at time $t\in\mathcal{T}$ when the price of the underlying at time $t$ is $s_{t}(j)$ and the element $\gamma_{t}(k)$ for $k\in\mathcal{I}_{t}$ provides the position in the underlying held between time $t\in[T-1]$ and time $t+1$, respectively. For the case of continuous marginal constraints, the corresponding duality was established by [22] for problems with a trivial initial marginal and a general terminal marginal in a continuous-time setting and by [16] for problems written on multiple assets with several given marginals in a discrete-time setting. See also [8, 15, 33, 44] for further work along these lines; see in particular [6, 54] for problems with multiple given marginals.
The dual (24) of the entropy-regularised problem (20) has no constraints — the subhedging constraint (27b) of problem (27) is incorporated in the objective function of problem (24). It follows from Proposition 4.1 and strong duality that the value of problem (24) converges to the value of problem (27) as the regularisation parameter $\varepsilon$ vanishes, however, showing convergence of optimal dual variables is less straightforward. It is our belief that an optimal solution of problem (24) serves as an approximation of an optimal subheding strategy of problem (27) for $\varepsilon>0$ small, but the matter should be investigated further.